# Mathematical Language Models: A Survey

### v3

@@ 5.1.3. Theorem Proving
Current research delves into exploring mathematical theorem datasets (e.g., INT (Wu et al., 2021), Feit-Thompson (Huang et al., 2019), and IsarStep (Li et al., 2021)) to develop novel machine learning-based theorem-proving strategies. For instance, the format of IsarStep is shown in 14. Huang et al. (Huang et al., 2019) introduced the Feit-Thompson dataset, encompassing 1,602 lemmas, expanding into 83,478 proof states for the Feit-Thompson theorem (Gonthier et al., 2013). IsarStep (Li et al., 2021) constitutes a non-synthetic dataset extracted from the largest repository of proofs handwritten by human experts in a theorem prover. Furthermore, NaturalProofs (Welleck et al., 2021) comprises 32,000 theorem statements and proofs, 14,000 definitions, and 2,000 other types of pages (e.g., axioms, corollaries). This dataset draws from three domains: broad-coverage data sourced from ProofWiki, deep-coverage data from the Stacks project, and low-resource data extracted from mathematics textbooks.
Figure 14. An example of IsarStep, taken from (Li et al., 2021).
Several formal mathematical libraries (Megill and Wheeler, 2019; Rudnicki, 1992; Grabowski et al., 2015) encompass a range of formal languages tailored for theorem proving (TP). Examples include MML (Grabowski et al., 2015), Coq (Bertot and Castéran, 2013), Lean (Han et al., 2022), Isabelle (Wenzel et al., 2008), and Hol light (Slind and Norrish, 2008). For instance, MML (Grabowski et al., 2015), also known as the Mizar Mathematical Library, was established to systematically construct a centralized, reusable knowledge base reflecting the standard foundations of mathematics, namely classical first-order logic and set theory. Yang et al. (Yang and Deng, 2019) compiled the CoqGym dataset, comprising 71,000 human-written proofs from 123 projects developed using the Coq proof assistant. LeanStep (Han et al., 2022) serves as the tactic proof dataset for the Lean theorem prover, featuring high-level human-written tactics alongside kernel-level proof terms. LISA (Jiang et al., 2021), one of the largest proof corpora for interactive theorem provers, contains 183,000 theorems and 2.16 million proof steps. This dataset is extracted from the Archive of Formal Proofs and the standard library of Isabelle. Kaliszyk et al. (Kaliszyk et al., 2017) introduced the HolStep dataset rooted in Higher-Order Logic (HOL) proofs. HOList (Bansal et al., 2019) includes almost 30,000 theorems and proofs across three corpora: core, complex, and flyspeck. The core corpus contains fundamental theorems necessary for defining tactics, while the complex corpus consists of theorems related to complex calculus. Flyspeck contains the majority of the lemmas and theorems related to the Kepler conjecture (Hales et al., 2017).
@@ 5.2. Benchmark Datasets
Various benchmark datasets have been introduced to assess the performance of different mathematical algorithms, such as MAWPS (Koncel-Kedziorski et al., 2016). Huang et al. (Huang et al., 2016) developed the Dolphin18K dataset, a large-scale and diverse collection containing over 18,000 annotated math word problems. The Mathematics dataset (Saxton et al., 2019) encompasses problems spanning arithmetic, algebra, probability, and calculus. These problems involve sequential questions and answers presented in a free-form textual input/output format. For assessing arithmetic understanding, Mishra et al. (Mishra et al., 2022b) introduced NumGLUE, a multi-task benchmark that evaluates AI systems across eight distinct tasks. miniF2F (Zheng et al., 2023) is a dataset designed to create a unified cross-system benchmark for neural theorem proving. It comprises 488 formal Olympiad-level mathematics problem statements, encompassing Metamath, Lean, Isabelle, and HOL Light.
Several datasets consider varying difficulty levels across educational stages, ranging from elementary school (Miao et al., 2020; Shi et al., 2022) to high school (Hendrycks et al., 2021b). Miao et al. (Miao et al., 2020) introduced the ASDiv corpus, comprising 2,305 math word problems (MWPs) categorized by problem type and elementary school grade level. The Multilingual Grade School Math Benchmark (MGSM) (Shi et al., 2022) leverages problems from GSM8K, translating them into 10 languages with the assistance of human annotators. The MATH dataset (Hendrycks et al., 2021b) contains 12,500 problems sourced from high school math competitions. Each problem within MATH is equipped with a step-by-step solution, enabling models to learn to generate answer derivations and explanations.
Furthermore, several datasets serve to evaluate foundation models (Zhong et al., 2023; Yuan et al., 2023b). AGIEval (Zhong et al., 2023) functions as a human-centric benchmark tailored to assess foundation models’ general abilities in tasks related to human cognition and problem-solving. This dataset is collected from standardized human-centric exams like college entrance exams, law school admission tests, math competitions, and lawyer qualification tests. Yuan et al. (Yuan et al., 2023b) introduced the arithmetic dataset MATH 401, designed to test the latest large language models such as GPT-4, ChatGPT, InstrctGPT, Galactica, and LLaMA. It features various arithmetic expressions and offers a comprehensive analysis of the capabilities of large language models.
@@ 5.3. Augmented Datasets
Augmented datasets serve to enhance existing datasets by incorporating additional samples or information. Roy et al. (Roy and Roth, 2018) introduced Aggregate by augmenting the AllArith dataset with 661 word problems from Perturb. Perturb contains problems that were manually pruned either for not yielding the desired solution $a$ or being too dissimilar from the input problem $p$. Large-scale language models like ChatGPT and GPT-4 have been utilized to expand datasets. For instance, MetaMathQA (Yu et al., 2023b) rephrases questions from multiple perspectives without introducing extra knowledge to bootstrap mathematical questions. The Math5K dataset (Li et al., 2023a) comprises 50,000 problem-solution pairs generated using GPT-4. These datasets leverage advanced language models to enrich the existing collection of problems and solutions.
@@
Several studies have expanded existing datasets by including supplementary information, such as code programs (Austin et al., 2021; Mishra et al., 2022a) and explanations (Kim et al., 2022). MathQA-Python (Austin et al., 2021) is a Python adaptation of the MathQA benchmark, comprising 23,914 problems. This dataset aims to evaluate models’ capabilities in synthesizing code from intricate textual descriptions. Lila (Mishra et al., 2022a) is a comprehensive mathematical reasoning benchmark composed of 23 diverse tasks. It extends 20 existing datasets by gathering task instructions and solutions as Python programs, offering not only correct answers but also explainable solutions. PEN (Kim et al., 2022) focuses on providing plausible and accurate explanations for solving algebraic word problems present in three benchmark datasets: ALG514, DRAW-1K, and MAWPS. Moreover, some studies have contributed statements and proofs based on existing datasets for theorem proving (Jiang et al., 2023b; Welleck et al., 2022). For example, miniF2F+informal (Jiang et al., 2023b) is a dataset consisting of manually curated informal statements and proofs aligned with formal statements from the miniF2F dataset. NaturalProofs-Gen (Welleck et al., 2022) adapts data from NATURALPROOFS, including theorem statements, proofs, definitions, and additional pages sourced from ProofWiki. Each proof follows a multi-step structure and references a variable number of sources.

### v4

@@ 5.1.3. Theorem Proving
Current research delves into exploring mathematical theorem datasets (e.g., INT (Wu et al., 2021), Feit-Thompson (Huang et al., 2019), and IsarStep (Li et al., 2021)) to develop novel machine learning-based theorem-proving strategies. Huang et al. (Huang et al., 2019) introduced the Feit-Thompson dataset, encompassing 1,602 lemmas, expanding into 83,478 proof states for the Feit-Thompson theorem (Gonthier et al., 2013). IsarStep (Li et al., 2021) constitutes a non-synthetic dataset extracted from the largest repository of proofs handwritten by human experts in a theorem prover. Furthermore, NaturalProofs (Welleck et al., 2021) comprises 32,000 theorem statements and proofs, 14,000 definitions, and 2,000 other types of pages (e.g., axioms, corollaries). This dataset draws from three domains: broad-coverage data sourced from ProofWiki, deep-coverage data from the Stacks, and low-resource data extracted from mathematics textbooks. Several formal mathematical libraries (Megill and Wheeler, 2019; Rudnicki, 1992; Grabowski et al., 2015) encompass a range of formal languages tailored for theorem proving (TP). Examples include MML (Grabowski et al., 2015), Coq (Bertot and Castéran, 2013), Lean (Han et al., 2022), Isabelle (Wenzel et al., 2008), and Hol light (Slind and Norrish, 2008). For instance, MML (Grabowski et al., 2015), also known as the Mizar Mathematical Library, was established to systematically construct a centralized, reusable knowledge base reflecting the standard foundations of mathematics, namely classical first-order logic and set theory. Yang et al. (Yang and Deng, 2019) compiled the CoqGym dataset, comprising 71,000 human-written proofs from 123 projects developed using the Coq proof assistant. LeanStep (Han et al., 2022) serves as the tactic proof dataset for the Lean theorem prover, featuring high-level human-written tactics alongside kernel-level proof terms. LISA (Jiang et al., 2021), one of the largest proof corpora for interactive theorem provers, contains 183,000 theorems and 2.16 million proof steps. This dataset is extracted from the Archive of Formal Proofs and the standard library of Isabelle. Kaliszyk et al. (Kaliszyk et al., 2017) introduced the HolStep dataset rooted in Higher-Order Logic (HOL) proofs. HOList (Bansal et al., 2019) includes almost 30,000 theorems and proofs across three corpora: core, complex, and flyspeck. The core corpus contains fundamental theorems necessary for defining tactics, while the complex corpus consists of theorems related to complex calculus. Flyspeck contains the majority of the lemmas and theorems related to the Kepler conjecture (Hales et al., 2017).
@@ 5.2. Benchmark Datasets
Various benchmark datasets have been introduced to assess the performance of different mathematical algorithms, such as MAWPS (Koncel-Kedziorski et al., 2016). Huang et al. (Huang et al., 2016) developed Dolphin18K, a diverse collection containing over 18,000 annotated MWPs. Mathematics (Saxton et al., 2019) encompasses problems spanning arithmetic, algebra, probability, and calculus. These problems involve sequential questions and answers presented in a free-form textual input/output format. For assessing arithmetic understanding, Mishra et al. (Mishra et al., 2022b) introduced NumGLUE, a multi-task benchmark that evaluates AI systems across eight distinct tasks. miniF2F (Zheng et al., 2023) is a dataset designed to create a unified cross-system benchmark for neural theorem proving. It comprises 488 formal Olympiad-level mathematics problem statements, encompassing Metamath, Lean, Isabelle, and HOL Light. FOLIO (Han et al., 2024) is a human-annotated, logically complex and diverse dataset for reasoning in natural language (NL), equipped with 1,430 first-order logic (FOL) annotations. Several datasets consider varying difficulty levels across educational stages, ranging from elementary to high school (Miao et al., 2020; Shi et al., 2022; Hendrycks et al., 2021b). Miao et al. (Miao et al., 2020) introduced ASDiv, comprising 2,305 MWPs categorized by problem type and elementary school grade level. The Multilingual Grade School Math Benchmark (MGSM) (Shi et al., 2022) leverages problems from GSM8K, translating them into 10 languages with the assistance of human annotators. The MATH dataset (Hendrycks et al., 2021b) contains 12,500 problems sourced from high school math competitions, where each problem within MATH is equipped with a step-by-step solution. Furthermore, several datasets serve to evaluate foundation models (Zhong et al., 2023; Yuan et al., 2023b). AGIEval (Zhong et al., 2023) functions as a human-centric benchmark tailored to assess foundation models’ general abilities in tasks related to human cognition and problem-solving. Yuan et al. (Yuan et al., 2023b) introduced the arithmetic dataset MATH 401, designed to test the latest large language models. With the development of multimodal LLMs (MLLMs), benchmarks for multimodal mathematical reasoning are proposed (Yue et al., 2024; Qiao et al., 2024; Chernyshev et al., 2024; Chen et al., 2024b; Zhang et al., 2025; Wang et al., 2024b). For example, MathVista (Lu et al., 2023a) is a popular English mathematical benchmark for MLLMs. Liu et al. (Liu et al., 2024) released a Chinese multimodal mathematical reasoning dataset to enhance and measure the performance of Multimodal LLMs. Moreover, MMMU-Math (Yue et al., 2024) and U-Math (Chernyshev et al., 2024) collect university-level problems to further evaluate the performance.
@@ 5.3. Augmented Datasets
Augmented datasets serve to enhance existing datasets by incorporating additional samples or information. Roy et al. (Roy and Roth, 2018) introduced Aggregate by augmenting the AllArith dataset with 661 word problems from Perturb. Perturb contains problems that were manually pruned either for not yielding the desired solution $a$ or being too dissimilar from the input problem $p$. Large-scale language models like ChatGPT and GPT-4 have been utilized to expand datasets. For instance, MetaMathQA (Yu et al., 2023b) rephrases questions from multiple perspectives without introducing extra knowledge to bootstrap mathematical questions. The Math5K dataset (Li et al., 2023a) comprises 50,000 problem-solution pairs generated using GPT-4. These datasets leverage advanced language models to enrich the existing collection of problems and solutions. Several studies have expanded existing datasets by including supplementary information, such as code programs (Austin et al., 2021; Mishra et al., 2022a) and explanations (Kim et al., 2022). MathQA-Python (Austin et al., 2021) is a Python adaptation of the MathQA benchmark, comprising 23,914 problems. This dataset aims to evaluate models’ capabilities in synthesizing code from intricate textual descriptions. Lila (Mishra et al., 2022a) is a comprehensive mathematical reasoning benchmark composed of 23 diverse tasks. It extends 20 existing datasets by gathering task instructions and solutions as Python programs, offering not only correct answers but also explainable solutions. PEN (Kim et al., 2022) focuses on providing plausible and accurate explanations for solving algebraic word problems present in three benchmark datasets: ALG514, DRAW-1K, and MAWPS. Moreover, some studies have contributed statements and proofs based on existing datasets for theorem proving (Jiang et al., 2023b; Welleck et al., 2022). For example, miniF2F+informal (Jiang et al., 2023b) is a dataset consisting of manually curated informal statements and proofs aligned with formal statements from the miniF2F dataset. NaturalProofs-Gen (Welleck et al., 2022) adapts data from NATURALPROOFS, including theorem statements, proofs, definitions, and additional pages sourced from ProofWiki. Each proof follows a multi-step structure and references a variable number of sources.
Table 2. Comparison of typical mathematical LMs. Pre-T: Pre-Training, FT: Finetuning, MM: Multimodal, Symbolic-S: Symbolic solvers, F-CoT: Foundational CoT, A-CoT: Advanced CoT, RL: Reinforce Learning
[…]
GPT-$f$ (Polu and Sutskever, 2020a) - Pre-T Symbolic-S
LISA (Jiang et al., 2021) - Pre-T Symbolic-S
MATH-PLM (Hendrycks et al., 2021b) GPT-2, GPT-3 Pre-T F-CoT
Minerva (Lewkowycz et al., 2022) PaLM Pre-T
MWP-BERT (Liang et al., 2022) BERT Pre-T
ControlMath (Chen et al., 2024c) LLaMA-2, Mistral FT F-CoT
EURUS (Yuan et al., 2024) Mistral, CodeLLaMA FT F-CoT DPO, KTO, NCA
MathGLM (Yang et al., 2023) GLM FT
WizardLM (Xu et al., 2023) LLaMA FT
Wizardmath (Luo et al., 2023) LLaMA-2 FT PPO
PaLM 2-L-Math (Liu et al., 2023b) PaLM-2-L FT F-CoT
Codex-math (Drori et al., 2022) Codex Prompt Program
Toolformer (Schick et al., 2023) GPT-J FT Calculator
LeanReasoner (Jiang et al., 2024) CodeLLaMA FT Symbolic-S
LINC (Olausson et al., 2023) StarCoder+ Prompt Symbolic-S F-CoT
MuMath-Code (Yin et al., 2024) Llama-2 FT Program
MAmmoTH (Yue et al., 2023) - FT Program F-CoT
InternLM-Math (Ying et al., 2024) InternLM2 Pre-T, FT Program F-CoT PPO
Qwen-Math (Yang et al., 2024) - Pre-T, FT Program F-CoT PPO
PAL (Gao et al., 2023) - Prompt Program F-CoT
CoT (Wei et al., 2022) - Prompt F-CoT
PromptPG-CoT (Lu et al., 2023c) - Prompt
BoostedPrompt (Pitis et al., 2023) - Prompt F-CoT
ToT (Yao et al., 2023b) - Prompt A-CoT
PoT (Chen et al., 2022a) - Prompt Program F-CoT
Self-check (Miao et al., 2023) - Prompt A-CoT
Self-Consistency (Wang et al., 2023d) - Prompt A-CoT
Diversity-of-Thought (Naik et al., 2023) - Prompt A-CoT
RAP (Hao et al., 2023a) - Prompt A-CoT
LATS (Zhou et al., 2023e) - Prompt A-CoT
LLM+P (Liu et al., 2023a) - Prompt Symbolic-S A-CoT
LLM+DP (Dagan et al., 2023) - Prompt Symbolic-S A-CoT
ISR-LLM (Zhou et al., 2023b) - Prompt A-CoT
MCR (Yoran et al., 2023) - Prompt A-CoT
Rank-verifier (Cobbe et al., 2021) - Prompt A-CoT
GRACE (Khalifa et al., 2023) - Prompt A-CoT
Reflexion (Shinn et al., 2023) - Prompt A-CoT
CR (Zhang et al., 2023c) - Prompt A-CoT
Step-Plan (Zhang et al., 2023b) - Prompt A-CoT
STaR (Zelikman et al., 2022) GPT-J FT A-CoT
V-STaR (Hosseini et al., 2024) LLaMA2, CodeLLaMA FT A-CoT DPO
Quiet-STaR (Zelikman et al., 2024) Mistral Pre-T A-CoT
REFT (Luong et al., 2024) Galactica, CodeLLaMA FT A-CoT PPO
SCoRe (Kumar et al., 2024) Gemini FT A-CoT SCoRe
HGS-PRM (Ma et al., 2023) LLaMA2, WizardMath FT A-CoT PPO
MCTSr (Zhang et al., 2024b) LLaMa-3 FT A-CoT
CoRe (Zhu et al., 2022) GPT-J, DeBERTa-large FT A-CoT
Marco-o1 (Zhao et al., 2024c) Qwen2 FT A-CoT
Step-DPO (Lai et al., 2024) Qwen2 FT A-CoT DPO
Flow-DPO (Deng and Mineiro, 2024) LLaMA3, Phi-3 FT A-CoT DPO
OmegaPRM (Luo et al., 2024) Gemini Pro, Gemma2 FT A-CoT PRM
SocraticLLM (Ding et al., 2024) Qwen2 FT A-CoT
MATHDIAL (Macina et al., 2023) T5 FT A-CoT
QwQ (Team, 2024) Qwen2.5 Pre-T Long CoT
o1 (OpenAI, 2024) - Pre-T Long CoT
GPT-4o (Hurst et al., 2024) - Pre-T A-CoT
GPT-4V (OpenAI, 2023) - Pre-T A-CoT
Qwen2-VL (Wang et al., 2024a) Qwen2-VL Pre-T F-CoT
LLaVA-o1 (Xu et al., 2024b) LLaVA FT Long CoT
Gemini (Team et al., 2023) Gemini Pre-T A-CoT
GLM-4V (GLM et al., 2024) ChatGLM Pre-T A-CoT
AtomThink (Xiang et al., 2024) LLaVA, EMOVA FT Long CoT PRM
Math-LLaVA (Shi et al., 2024) LLaVA FT F-CoT
M-STAR (Anonymous, 2024) MiniCPM-V FT A-CoT PRM
UnAC (wang and Fu, 2024) GPT-4V, Gemini, LLaVA Prompt F-CoT
[…]
Table 3. Performance of various mathematical LMs in terms of accuracy.
[…]
MATH-PLM (Hendrycks et al., 2021b) 6.9 - - - - - - - -
Minerva (Lewkowycz et al., 2022) 50.3 78.5 75.0 - - - - - -
MWP-BERT (Liang et al., 2022) - - - - - 76.6 - - -
ControlMath (Chen et al., 2024c) 9.5 49.3 - - - - 51.6 - -
EURUS (Yuan et al., 2024) 41.7 62.8 - 93.0 - - 90.4 - -
Wizardmath (Luo et al., 2023) 22.7 81.6 - - - - - -
Codex-math (Drori et al., 2022) 81.8 - - - - - - - -
Toolformer (Schick et al., 2023) - - - 40.4 - - 29.4 - -
LeanReasoner (Jiang et al., 2024) - - - - - - - 82.6 -
LINC (Olausson et al., 2023) - - - - - - - 72.5 -
MuMath-Code (Yin et al., 2024) 55.1 90.7 - - - - - - -
MAmmoTH (Yue et al., 2023) 44.2 76.7 - - - - - - 61.4
InternLM-Math (Ying et al., 2024) 37.7 82.6 - - - - - - -
Qwen-Math (Yang et al., 2024) 66.8 91.6 82.8 - 63.3 86.3 - - -
PAL (Gao et al., 2023) - 72.0 - 79.6 - - 79.4 - -
BoostedPrompt (Pitis et al., 2023) - 87.1 - - - - - - 63.8
PoT (Chen et al., 2022a) - 80.0 - - - 89.1 - - 58.6
Self-check (Miao et al., 2023) 51.3 88.1 - - 81.2 - - - -
Self-Consistency (Wang et al., 2023d) - 78.0 - 87.8 - - 86.8 - 52.0
Diversity-of-Thought (Naik et al., 2023) - 96.3 - - - - - - 81.7
GRACE (Khalifa et al., 2023) - 36.3 - - - 84.4 68.6 - -
CR (Zhang et al., 2023c) 72.2 - - - - - - 98.0 -
REFT (Luong et al., 2024) - 75.3 - - - 71.8 79.2 - -
SCoRe (Kumar et al., 2024) 64.4 - - - - - - - -
HGS-PRM (Ma et al., 2023) 13.7 65.4 - - - - - - -
MCTSr (Zhang et al., 2024b) 58.2 96.6 - - 11.8 - - - -
CoRe (Zhu et al., 2022) - 63.2 - 90.5 - - - - -
Step-DPO (Lai et al., 2024) 70.8 94.0 - - - - - - -
Flow-DPO (Deng and Mineiro, 2024) 38.6 71.3 - - - - - - -
OmegapRM (Luo et al., 2024) 69.4 93.6 - - - - - - -
QwQ (Team, 2024) 90.6 - - - 50.0 - - - -
GLM-4-9B-Chat (GLM et al., 2024) 50.6 79.6 72.4 - - - - - -
[…]
o1 (OpenAI, 2024) 94.8 - 92.3 - 83.3 78.1 73.9 - -
GPT-4o (Hurst et al., 2024) 60.3 - 88.0 - 13.4 69.2 63.8 29.02 30.4
GPT-4V (OpenAI, 2023) - - - - - 56.8 49.9 - 22.8
Qwen2-VL (Wang et al., 2024a) - - 46.2 - - 64.5 70.5 43.0 25.9
LLaVA-o1 (Xu et al., 2024b) 64.0 - - - - - 54.8 - -
Gemini (Team et al., 2023) 67.7 - 85.9 - - 62.2 63.9 41.9 17.7
GLM-4V (GLM et al., 2024) - - - - - 47.2 - - -
AtomThink (Xiang et al., 2024) - - - - - - 53.3 - 40.5
Math-LLaVA (Shi et al., 2024) - - - - 38.3 38.3 46.6 - 15.7
M-STAR (Anonymous, 2024) - - - - - - 59.5 - -
UnAC (wang and Fu, 2024) - - - - - 59.7 56.6 - -
[…]
6. Analysis and Discussion
To gain a deeper understanding of the advantages of typical mathematical LMs, we conduct a comprehensive comparison of their characteristics and performance (Table 2 and Table 3). Table 2 provides a detailed overview of mathematical LLMs, their architectures, training strategies, reasoning capabilities, and other advanced features. From the table, we observe the following findings. First, CoT has become a critical enabler for mathematical reasoning. Additionally, with the introduction of OpenAI’s o1 model, an increasing number of studies are focusing on Long CoT. Second, only a few models (e.g., LLaVA, M-STAR) support multimodal input (text + image). This indicates that most mathematical LLMs focus solely on text-based reasoning. As tasks increasingly require interpreting diagrams and visual data, multimodal models could play a more prominent role in future developments. Third, models like Toolformer and LeanReasoner integrate external tools (e.g., calculators or symbolic solvers). These tools complement the LLMs’ reasoning by handling computations or symbolic operations that the models alone might struggle with. Fourth, reinforcement learning (RL) is primarily used to refine reasoning strategies and align model outputs with user expectations or task requirements. However, only a small subset of models uses RL techniques, such as Direct Preference Optimization (DPO) (Rafailov et al., 2024) or Proximal Policy Optimization (PPO) (Schulman et al., 2017), due to the instability of RL. Fifth, many recent models (e.g., CoT, BoostedPrompt, PoT) rely on prompt engineering to tackle complex problems without requiring heavy computational resources for fine-tuning. Fine-tuning is essential for most advanced mathematical models (e.g., WizardMath, EURUS) to enhance mathematical reasoning capabilities. Table 3 compares the performance of various mathematical LMs across different datasets. The methods can be classified into two categories: Text-based Methods and Multi-modal Methods. Such benchmarks are widely recognized and serve to assess mathematical capabilities in diverse contexts. Here are some conclusions based on the data: First, o1 stands out as the most accurate model across most datasets, achieving the highest scores in MATH (94.8%), MMLU (92.3%), AIME (83.3%), MMMU (78.1%) and MathVista (73.9%). It indicates that slow thinking with long CoT improves the performance of mathematical reasoning effectively. Second, methods like EURIUS and PoT show dataset-specific strengths, suggesting that tailored approaches can yield better results for certain problem types (e.g., formal proofs, complex reasoning). For example, CR achieves the best performance on the FOLIO Dataset with 98.0%, highlighting its capability for handling formal logic problems. Third, some Multi-modal benchmarks, such as CMM-Math (Liu et al., 2024) and Math-V (Wang et al., 2024b), show relatively low accuracy across all methods, indicating a need for further research and improvement in these areas.